# A Note on Sample Complexity of Interactive Imitation Learning with Log Loss

### v1

@@ B.1 Decoupling State and Action Sequences by Decoupled Hellinger Distance
In this section, we demonstrate that similar to $D^{2}_{\text{H}}\left(\mathbb{P}^{\pi}(s_{1:H},a_{1:H}),\mathbb{P}^{\pi^{\mathrm{E}}}(s_{1:H},a_{1:H})\right)$ [4], the decoupled Hellinger distance $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right]$ that decouples states and actions is also proportionally lower bounded by a constant factor of $\rho(\pi\parallel\pi^{\mathrm{E}})$. The following two lemma shows that such relationship holds for both Markovian policies and their first-step mixtures.
@@
$\frac{1}{2}\cdot\rho(\pi\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
@@
Let $\pi^{\mathrm{E}}$ be a deterministic policy, and let $\pi_{u}$ be a first-step mixture of Markovian policies. Then we have
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
To prove these two lemmas, we first prove a special case, i.e. Lemma 11 with first-step mixture of deterministic policies. To facilitate the proofs, we introduce the following additional notations:
@@
Let $\pi^{\mathrm{E}}$ be a deterministic Markovian policy, and let $\pi_{u}$ be a first-step mixture of deterministic Markovian policies (elements of $\mathcal{B}^{\mathrm{Det}}$). Then we have that
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
The key idea in the following proof is to lower bound $\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right],$ which reflects the asymmetric roles of the two appearances of $\pi_{u}$’s, using a symmetric formulation via function $F$ (as shown in (6)).
@@
Recall the first-step mixture policy in Definition 1, we start by rewriting
@@
Next, we turn to analyzing $D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))$. Since the deterministic expert induces a delta mass distribution over actions, we apply the elementary fact about the Hellinger distance with delta mass distribution stated in Lemma 18, yielding:
$\frac{1}{2}\parallel\pi_{u}(s_{1:H})-\pi^{\mathrm{E}}(s_{1:H})\parallel_{1}\leq D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H})).$
@@
$1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H}))\leq D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H})).$ (4)
@@
$\sum_{s_{1:H}}\mathbb{P}^{\pi_{u}}(s_{1:H})(1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H})))\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$ (5)
@@
Since $\pi_{u}$ is a first-step mixture of policies in $\mathcal{B}^{\mathrm{Det}}$ with weight $u$, we have $\mathbb{P}^{\pi_{u}}(s_{1:H})=\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\mathbb{P}^{\nu}(s_{1:H})$. This allows us to rewrite ()(*)( ∗ ) using the definition of $F(\nu;\nu^{\prime},\pi^{\mathrm{E}})$ as:
()(*)=( ∗ ) = $\sum_{s_{1:H}}\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\mathbb{P}^{\nu}(s_{1:H})\sum_{\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu^{\prime})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]$ (6)
@@
((*$\geq$( ∗ ) ≥ $\frac{1}{2}\cdot\frac{1}{2}\cdot\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})\left(\rho(\nu\parallel\pi^{\mathrm{E}})+\rho(\nu^{\prime}\parallel\pi^{\mathrm{E}})\right)$
@@
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq(*)\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
@@
In words, the four subsets divide state sequences into cases where: (1) both $\nu,\nu^{\prime}$ agree with the $\pi^{\mathrm{E}}$ throughout, (2)&\&&(3) one of $\nu,\nu^{\prime}$ makes its first mistake earlier than the other, and (4) $\nu,\nu^{\prime}$ make their first mistake at the same time. It can now be easily seen that each $s_{1:H}\in{\mathcal{S}}^{H}$ lies in exactly one of such $\mathcal{X}_{i}$, and
@@
$\mathcal{X}_{2}^{h}:=$$\{s_{1:H}\;|\;\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\};$ (8)
$\mathcal{X}_{3}^{h}:=$$\{s_{1:H}\;|\;\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\};$
$\mathcal{X}_{4}^{h}:=$$\{s_{1:H}\;|\;\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\}.$
@@
$\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu}(s_{1:h})=$ $\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}P_{0}(s_{1})\prod_{h^{\prime}=1}^{h-1}P_{h}(s_{h+1}|s_{h},\nu_{h}(s_{h}))$ (9)
$=$ $\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}P_{0}(s_{1})\prod_{h^{\prime}=1}^{h-1}P_{h}(s_{h+1}|s_{h},\nu^{\prime}_{h}(s_{h}))=\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu^{\prime}}(s_{1:h}).$
@@
We conclude the proof by plugging (15) into (14).
B.1.1 Proof of Lemma 9
@@
The key to this proof is showing that any Markovian policy $\pi$ is equivalent, in terms of action distribution on any state sequence, to a first-step mixture of a set of deterministic Markovian policies. This leads to equivalence on trajectory distribution and decoupled Hellinger distance. To clarify further, we present the following claim, which allows us to apply guarantees for mixtures of deterministic policies in Lemma 11.
@@
For a Markovian policy $\pi$, there exists a first-step mixture of deterministic policy $\pi_{u}$ such that for any $s_{1:H}\in{\mathcal{S}}^{H}$, 1. $\pi(s_{1:H})=\pi_{u}(s_{1:H})$, and 2. $\mathbb{P}^{\pi}(s_{1:H})=\mathbb{P}^{\pi_{u}}(s_{1:H})$.
@@
We now verify the first item. By first-step mixture, we rewrite $\pi_{u}(a_{1:H}\parallel s_{1:H})$ as
@@
$\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right]=\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
@@ B.1.2 Proof of Lemma 10
By Claim 13, any Markovian policy can be viewed as a first-step mixture of $A^{SH}$ deterministic policies from $\mathcal{B}^{\mathrm{Det}}$, then any first-step mixture of Markovian policies $\pi_{u}$ can also be viewed as a first-step mixture of $A^{SH}$ deterministic policies from $\mathcal{B}^{\mathrm{Det}}$. The proof follows by applying Lemma 11. ∎
B.2 New Guarantees for DAgger with Trajectory-wise Annotation
@@
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
In the following, we first demonstrate that the performance difference between expert and the the uniform first-step mixture of any Markovian policy sequence $\{\pi^{n}\}_{n=1}^{N}$ is upper-bounded by $2R\operatorname{OnEst}^{\mathrm{Traj}}_{N}/N$, and then show the trajectory-wise sample complexity of Algorithm 2 in Theorem 4.
@@
For any MDP $\mathcal{M}$, deterministic expert $\pi^{\mathrm{E}}$, and sequence of policies $\left\{\pi^{n}\right\}_{n=1}^{N}$, each of which can be Markovian or a first-step mixture of Markovian policies, their first step uniform mixture policy $\hat{\pi}$ satisfies.
@@
By Lemma 9 and Lemma 10, for each $\pi^{n}$, whether it is Markovian or a first-step mixture of Markovian policies, the following holds:
$\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right]\geq\frac{1}{2}\rho(\pi^{n}\parallel\pi^{\mathrm{E}}).$
@@
$\frac{\operatorname{OnEst}^{\mathrm{Traj}}_{N}}{N}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(s^{n}_{1:H}),\pi^{\mathrm{E}}(s^{n}_{1:H}))\right]\geq\frac{1}{2N}\sum_{n=1}^{N}\rho(\pi^{n}\parallel\pi^{\mathrm{E}})=\frac{1}{2}\rho(\hat{\pi}\parallel\pi^{\mathrm{E}}),$
where we apply the fact that $\hat{\pi}$ is a first-step mixture of $\{\pi^{n}\}_{n=1}^{N}$ with uniform weights. Finally, we conclude the proof by applying Lemma 20. ∎
@@
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq 2R\frac{\log(B)+2\log(1/\delta)}{N}.$
@@
The proof closely follows Proposition C.2 in [4], tailored for another DAgger variant. However, in this case, we leverage the distribution of the state sequence $s_{1:H}$ instead of the per-step state distribution.
Observe that the log-loss functions passed through $\textbf{Alg}_{\text{Est}}$, $\ell^{n}(u)$ is of the form
$\ell^{n}(u)=\log(1/\pi_{u}(a^{n,*}_{1:H}\parallel s^{n}_{1:H}))=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a^{n,*}_{1:H}\parallel s^{n}_{1:H})}\right).$
It can be observed that $\ell^{n}$’s are 1-exp-concave. Therefore, implementing $\textbf{Alg}_{\text{Est}}$ using the exponential weights algorithm (Proposition 21) ensures that the following bound holds almost surely:

### v2

@@ E.3 Decoupling State and Action Sequences by Decoupled Hellinger Distance
In this section, we demonstrate that similar to $D^{2}_{\text{H}}\left(\mathbb{P}^{\pi}(s_{1:H},a_{1:H}),\mathbb{P}^{\pi^{\mathrm{E}}}(s_{1:H},a_{1:H})\right)$ [foster2024behavior], the decoupled Hellinger distance $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right]$ that decouples states and actions is also proportionally lower bounded by a constant factor of $\rho(\pi\parallel\pi^{\mathrm{E}})$. The following two lemma shows that such relationship holds for both Markovian policies and their first-step mixings.
@@
$\frac{1}{2}\cdot\rho(\pi\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
@@
Let $\pi^{\mathrm{E}}$ be a deterministic policy, and let $\pi_{u}$ be a first-step mixing of Markovian policies. Then we have
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
To prove these two lemmas, we first prove a special case, i.e. Lemma 30 with first-step mixing of deterministic policies. To facilitate the proofs, we introduce the following additional notations:
@@
Let $\pi^{\mathrm{E}}$ be a deterministic Markovian policy, and let $\pi_{u}$ be a first-step mixing of deterministic Markovian policies (elements of $\mathcal{B}^{\mathrm{Det}}$). Then we have that
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
The key idea in the following proof is to lower bound $\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right],$ which reflects the asymmetric roles of the two appearances of $\pi_{u}$’s, using a symmetric formulation via function $F$ (as shown in (9)).
@@
Recall the first-step mixing policy in Definition 24, we start by rewriting
@@
Next, we turn to analyzing $D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))$. Since the deterministic expert induces a delta mass distribution over actions, we apply the elementary fact about the Hellinger distance with delta mass distribution stated in Lemma 35, yielding:
$\frac{1}{2}\parallel\pi_{u}(s_{1:H})-\pi^{\mathrm{E}}(s_{1:H})\parallel_{1}\leq D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H})).$
@@
$1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H}))\leq D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H})).$ (7)
@@
$\sum_{s_{1:H}}\mathbb{P}^{\pi_{u}}(s_{1:H})(1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H})))\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$ (8)
@@
Since $\pi_{u}$ is a first-step mixing of policies in $\mathcal{B}^{\mathrm{Det}}$ with weight $u$, we have $\mathbb{P}^{\pi_{u}}(s_{1:H})=\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\mathbb{P}^{\nu}(s_{1:H})$. This allows us to rewrite ()(*) using the definition of $F(\nu;\nu^{\prime},\pi^{\mathrm{E}})$ as:
$(*)=$ $\sum_{s_{1:H}}\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\mathbb{P}^{\nu}(s_{1:H})\sum_{\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu^{\prime})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]$ (9)
@@
$(*)\geq$ $\frac{1}{2}\cdot\frac{1}{2}\cdot\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})\left(\rho(\nu\parallel\pi^{\mathrm{E}})+\rho(\nu^{\prime}\parallel\pi^{\mathrm{E}})\right)$
@@
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq(*)\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
@@
In words, the four subsets divide state sequences into cases where: (1) both $\nu,\nu^{\prime}$ agree with the $\pi^{\mathrm{E}}$ throughout, (2)&\&(3) one of $\nu,\nu^{\prime}$ makes its first mistake earlier than the other, and (4) $\nu,\nu^{\prime}$ make their first mistake at the same time. It can now be easily seen that each $s_{1:H}\in{\mathcal{S}}^{H}$ lies in exactly one of such $\mathcal{X}_{i}$, and
@@
$\mathcal{X}_{2}^{h}$= $\{s_{1:H}\;|\;\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\};$ (11)
$\mathcal{X}_{3}^{h}$= $\{s_{1:H}\;|\;\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\};$
$\mathcal{X}_{4}^{h}$= $\{s_{1:H}\;|\;\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\}.$
@@
$\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu}(s_{1:h})=$ $\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}P_{0}(\mathbf{E^{\prime}})\prod_{h^{\prime}=1}^{h-1}P_{h}(s_{h+1}|s_{h},\nu_{h}(s_{h}))$ (12)
$=$ $\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}P_{0}(\mathbf{E^{\prime}})\prod_{h^{\prime}=1}^{h-1}P_{h}(s_{h+1}|s_{h},\nu^{\prime}_{h}(s_{h}))=\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu^{\prime}}(s_{1:h}).$
@@
We conclude the proof by plugging (18) into (17). ∎
E.3.1 Proof of Lemma 28
@@
The key to this proof is showing that any Markovian policy $\pi$ is equivalent, in terms of action distribution on any state sequence, to a first-step mixing of a set of deterministic Markovian policies. This leads to equivalence on trajectory distribution and decoupled Hellinger distance. To clarify further, we present the following claim, which allows us to apply guarantees for mixtures of deterministic policies in Lemma 30.
@@
For a Markovian policy $\pi$, there exists a first-step mixing of deterministic policy $\pi_{u}$ such that for any $s_{1:H}\in{\mathcal{S}}^{H}$, 1. $\pi(s_{1:H})=\pi_{u}(s_{1:H})$, and 2. $\mathbb{P}^{\pi}(s_{1:H})=\mathbb{P}^{\pi_{u}}(s_{1:H})$.
@@
We now verify the first item. By first-step mixing, we rewrite $\pi_{u}(a_{1:H}\parallel s_{1:H})$ as
@@
$\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right]=\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
@@ E.3.2 Proof of Lemma 29
By Claim 32, any Markovian policy can be viewed as a first-step mixing of $A^{SH}$ deterministic policies from $\mathcal{B}^{\mathrm{Det}}$, then any first-step mixing of Markovian policies $\pi_{u}$ can also be viewed as a first-step mixing of $A^{SH}$ deterministic policies from $\mathcal{B}^{\mathrm{Det}}$. The proof follows by applying Lemma 30. ∎
E.4 New Guarantees for DAgger Variant with Trajectory-wise Annotation
@@
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
In the following, we first demonstrate that the performance difference between expert and the the uniform first-step mixing of any Markovian policy sequence $\{\pi^{n}\}_{n=1}^{N}$ is upper-bounded by $2R\operatorname{OnEst}^{\mathrm{Traj}}_{N}/N$, and then show the trajectory-wise sample complexity of Algorithm 3 in Theorem 27.
@@
For any MDP $\mathcal{M}$, deterministic expert $\pi^{\mathrm{E}}$, and sequence of policies $\left\{\pi^{n}\right\}_{n=1}^{N}$, each of which can be Markovian or a first-step mixing of Markovian policies, their first step uniform mixture policy $\hat{\pi}$ satisfies.
@@
By Lemma 28 and Lemma 29, for each $\pi^{n}$, whether it is Markovian or a first-step mixing of Markovian policies, the following holds:
$\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right]\geq\frac{1}{2}\rho(\pi^{n}\parallel\pi^{\mathrm{E}}).$
@@
$\frac{\operatorname{OnEst}^{\mathrm{Traj}}_{N}}{N}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s^{n}_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s^{n}_{1:H}))\right]\geq\frac{1}{2N}\sum_{n=1}^{N}\rho(\pi^{n}\parallel\pi^{\mathrm{E}})=\frac{1}{2}\rho(\hat{\pi}\parallel\pi^{\mathrm{E}}),$
where we apply the fact that $\hat{\pi}$ is a first-step mixing of $\{\pi^{n}\}_{n=1}^{N}$ with uniform weights. Finally, we conclude the proof by applying Lemma 37. ∎
@@
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq 2R\frac{\log(B)+2\log(1/\delta)}{N}.$
@@
The proof closely follows Proposition C.2 in [foster2024behavior], tailored for another DAgger variant. However, in this case, we leverage the distribution of the state sequence $s_{1:H}$ instead of the per-step state distribution.
Observe that the log loss functions passed through online learning oracle $\mathbb{A}$, $\ell^{n}(\pi)$ is of the form
$\ell^{n}(\pi)=\log(1/\pi_{u}(a^{n,*}_{1:H}\parallel s^{n}_{1:H}))=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a^{n,*}_{1:H}\parallel s^{n}_{1:H})}\right).$
It can be observed that $\ell^{n}$’s are 1-exp-concave. Therefore, implementing $\mathbb{A}$ using the exponential weights algorithm (Proposition 38) ensures that the following bound holds almost surely: